Water enters the condenser at 30 degrees Celsius and leaves at 60 degrees Celsius. What is the temperature difference in degrees Rankine


1
Expert's answer
2022-03-06T15:16:40-0500

The temperature difference in degrees Rankine

"\Delta T\:{^\circ\rm R}=\frac{9}{5}\Delta T\:{^\circ\rm C}"

"\Delta T\:{^\circ\rm R}=\frac{9}{5}*(60-30)\:{^\circ\rm C}=54\:{^\circ\rm R}"